# Canonical Policy: Learning Canonical 3D Representation for Equivariant Policy

### v1

@@ VI Real Robot Experiments
In this section, we evaluate canonical policy on two real-world platforms: the Franka Emika Panda and UR5. We describe the experimental setup, compare against both image-based and point cloud-based baselines, and assess generalization under appearance and viewpoint shifts, as well as data efficiency in low-data regimes.
@@ VI-A Real-World Experimental Setup
Franka Emika Panda: The first system consists of a Franka Emika Panda robot arm [52] using a relative control (see Section IV-C), equipped with a pair of Fin-Ray fingers [53] and a fixed Intel RealSense D415 [54] RGB-D camera. Human demonstrations are collected via a 6-DoF 3DConnexion SpaceMouse [55], with observations and actions recorded at 10 Hz. During policy rollout, we apply DDIM [56] sampling with 20 denoising steps for faster inference.
@@
[…]
@@
We compare our method with the representative point cloud baseline DP3 [3], and use CP-SO2 as our canonical policy variant due to the upright alignment of the observation $z$-axis. To assess performance across modalities, we additionally evaluate two state-of-the-art image-based baselines: Diffusion Policy (DP) [1] and the equivariant variant EquiDiff [9]. Input images are resized from $640\times 480$ to $256\times 256$, then center-cropped to $230\times 230$ before being fed into the encoder.
UR5: To test under different control and sensing conditions, we also conduct experiments on a UR5 [57] robot using absolute control (see Section IV-C). As shown in Fig. 11, the system is equipped with a WSG 50 gripper [58] for precise width control and Fin-Ray fingers for compliant grasping. A single RealSense D415 is mounted on a mobile workstation to provide RGB-D input. The mobility of the camera facilitates easy variation of viewpoints. Demonstrations are collected via the SpaceMouse. We follow the same point cloud generation and cropping procedure as described earlier.
On the UR5 platform, we conduct experiments to evaluate the data efficiency and robustness to camera viewpoint changes of canonical policy, comparing it against both image-based and point cloud-based policies.
@@ VI-B Benchmarking Across Diverse Tasks
TABLE V: Success rates over 10 trials for real-world tasks on the Franka Emika platform.
Modality Policy Block Stacking Shoe Alignment Can Insertion Table Organization
@@
CP-SO2 7/10 6/10 4/10 3/10
Table V summarizes the quantitative performance of different policies with various input modalities across four real-world tasks, evaluated over 10 trials on the Franka platform. CP-SO2 consistently achieves the highest success rates compared to other policies as task difficulty increases. For simpler tasks such as Block Stacking and Shoe Alignment, CP-SO2 achieves 70% and 60% success rates, outperforming the second-best policies by 40% and 10%, respectively. Besides, in these two tasks, policies based on point cloud modalities achieve higher success rates compared to those relying on image modalities.
@@
Fig. 12 illustrates the variability of policy rollouts over 10 trials. We sum the initial state observation of each trial for CP-SO2 and compute the average, which confirms that each trial differs from the others, ensuring sufficient randomness for fair evaluation. Therefore, the quantitative results reported in Table V validate the effectiveness of the proposed canonical representation, which outperforms baselines across both point cloud and image modalities in real-world scenarios.
@@ VI-C Generalization to Unseen Object Appearances
Figure 14: Visualization of randomized initial configurations under appearance changes.
@@
TABLE VI: Success rates over 10 trials under color variations on the Franka Emika platform.
Modality Policy Block Stacking Shoe Alignment Can Insertion Table Organization
@@
CP-SO2 6/10 6/10 4/10 2/10
TABLE VII: Success rates over 10 trials under both color and shape variations on the Franka Emika platform.
@@
CP-SO2 5/10 2/10
TABLE VIII: Success rates over 10 trials under varying numbers of demonstrations and shoe colors on the UR5 platform.
@@
CP-SO2 4/10 6/10 6/10 7/10
TABLE IX: Success rates over 10 trials under varying camera viewpoints and shoe colors on the UR5 platform.
@@
In this section, to thoroughly evaluate the generalizability of canonical policy, we modify the appearance of each task by altering object colors and shapes. The same trained policies for each task are then used to perform rollouts in the modified environments. The appearance changes are illustrated in Fig. 13.
@@
Specifically, in the Block Stacking task, both the I-shaped and T-shaped blocks are recolored to blue, representing a significant color shift from their original appearance. In the Shoe Alignment task, the shoes used during training are replaced with white shoes. Furthermore, two additional shoe types: leather shoes and hiking boots are introduced, differing substantially in both color and shape from the original shoes. In the Can Insertion task, the original black holder and orange can are replaced with a white holder and green can, respectively. Finally, in the Table Organization task, only the color of the small block is changed from black to white. Each new scenario is evaluated over 10 trials, and the randomness across trials is illustrated in Fig. 14.
@@
The experimental results under color variations are summarized in Table VI. Since point cloud-based policies are trained without color information, their performance under appearance changes is significantly more robust compared to image-based policies, as evidenced by the comparison with Table V. Notably, CP-SO2 consistently achieves the highest success rates across all tasks, outperforming the second-best policy by 30%, 30%, 10%, and 20% for tasks of increasing difficulty. Interestingly, Equidiff achieves the second-best performance in the Can Insertion task, outperforming DP3 by 20%, but shows zero success rates in other tasks, similar to DP. This phenomenon indicates that while Equidiff exhibits some degree of generalization, it remains inferior to canonical policy, as CP-SO2 maintains stable and high performance across all tasks under appearance changes.
@@
Table VII further assesses the generalization capability of each policy under simultaneous changes in both color and shape. CP-SO2 consistently achieves the highest success rates, outperforming the second-best policy by 20% in the Leather Shoe Alignment task. In the Hiking Boot Alignment task, DP3 achieves performance comparable to CP-SO2. This observation can be explained by the fact that, in the original training set, the shoes were low-top, leading the end-effector to learn grasping motions at relatively low heights. When encountering high-top hiking boots, the end-effector often follows a lower approach trajectory and collides with the boot during the grasping attempt, significantly increasing the difficulty of successful grasping. Under such challenging conditions, the performance of point cloud-based policies remains similar.
VI-D Data Efficiency and Robustness to Viewpoint Variations
@@
In addition to evaluating policy performance under nominal conditions, we assess the generalization capability of canonical policy in real-world settings with limited data and varying observation conditions. Specifically, we study two challenging scenarios on the UR5 platform: (i) data efficiency under varying numbers of demonstrations, and (ii) robustness to camera viewpoint changes.
@@
Table VIII summarizes the results under varying numbers of training demonstrations, comparing each method using 50 and 100 human demonstrations collected with black shoes. Each configuration is evaluated over 10 rollout trials. When trained with only 50 demonstrations, CP-SO2 is the only method that achieves non-zero success rates, reaching 40% on black shoes and 60% on white shoes. In contrast, all other baselines, including the image-based policies DP and Equidiff as well as the point cloud-based baseline DP3, fail completely under this low-data setting.
@@
As the number of demonstrations increases to 100, CP-SO2’s performance further improves to 60% and 70% on black and white shoes, respectively, outperforming the second-best method DP3 by 30% and 50%. Notably, while CP-SO2 and DP3 exhibit clear improvements with additional training data, the performance of image-based policies remains stagnant at 0%.
@@
As shown in Tables V and VIII, all baselines (DP, DP3, and EquiDiff) perform worse on the UR5 compared to the Franka Panda. We attribute this to differences in the low-level control setups. In the Franka Panda experiments, we utilized an Operational Space Controller, which helped smooth out the jerky actions produced by the learned policies. In contrast, the UR5 setup relied on high-precision pose tracking without such smoothing. As illustrated in our supplementary video, under a single-camera setup and with only 50 or 100 demonstrations, baseline policies frequently produced unstable motions on the UR5, leading to degraded task performance. These results further underscore the consistent and robust performance of canonical policy across different robotic platforms.
@@
An interesting observation is that CP-SO2 achieves higher success rates on white shoes, despite being trained exclusively on black shoes, underscoring the strong generalization capability of canonical policy. From a point cloud perspective, black and white shoes share the same geometric structure, enabling a policy trained on black shoes to naturally generalize to white ones. Nevertheless, the point clouds of white shoes introduce subtle variations compared to black shoes in the training set. We believe that this slight randomness in observations may actually enhance policy robustness, a phenomenon frequently observed in imitation learning, potentially explaining the improved performance on white shoes. Introducing controlled randomness into observations is a common strategy to improve robustness of policies, as seen in techniques such as random image cropping in Diffusion Policy [1] and color jitter or random rotations in UMI training [12].
@@
To evaluate the robustness of canonical policy to camera viewpoint shifts, we fix the number of training demonstrations to 50 and rotate the mobile workstation to simulate different viewing angles. As shown in Fig. 15, the leftmost two columns (Left-Angled View) correspond to the original camera position used during training. The middle columns (Frontal View) represent a moderate rotation of approximately $15^{\circ}$, where the camera is positioned in front of the workspace. The rightmost columns (Right-Angled View) reflect a more significant rotation of approximately $30^{\circ}$, inducing a larger deviation from the training viewpoint.
@@
A change in camera viewpoint is theoretically equivalent to an inverse transformation of the scene while keeping the camera fixed. This experimental setup therefore enables us to validate the theoretical foundation of our canonical representation. As the mobile workstation performs rigid-body translations and in-plane rotations, the $z$-axis remains consistently aligned upright, ensuring the applicability of CP-SO2, which is designed to be SO(2)-equivariant about the vertical axis.
@@
Table IX summarizes the success rates under different viewpoint shifts. When the viewpoint is shifted to the Frontal View, CP-SO2 maintains strong performance with success rates of 40% and 60% for black and white shoes, respectively. In contrast, all baseline policies, including the image-based DP and Equidiff and the point cloud-based DP3, fail completely under the same conditions.
When the camera is further rotated to the Right-Angled View, the performance of CP-SO2 drops to 30% for both shoe types. This degradation can be attributed to the single-camera setup, where larger viewpoint shifts result in significantly different observed point clouds. Since these new observations cover previously unseen regions, the policy may struggle to generalize to the resulting features. Conversely, in the small-angle case (Frontal View), the overlap between training and test point clouds remains high, allowing CP-SO2 to leverage its geometric equivariance and map transformed inputs to a consistent canonical pose, as discussed in Section IV-D. This enables improved generalization under viewpoint variations.

### v2

@@ 48 $85\!\pm\!3$ $87\!\pm\!4$ $76\!\pm\!5$ $83\!\pm\!6$
V-E Reliability Analysis of the Rotation Estimator
In Section III-C, we introduced the theoretical foundation of the canonical policy, based on the concept of equivariant groups. Section IV-A further proves that all elements within an equivariant group can be transformed into a shared canonical representation. However, this definition is highly restrictive: even a tiny deviation, such as a single point difference, may break the rigid-transformation constraint and cause point clouds to fall into different groups, which introduces more variation for the policy. In practice, neural networks exhibit robustness to such noise through parameter optimization.
We design an experiment to evaluate the reliability of the $\mathrm{SO}(3)$-equivariant network shown in Fig. 3 under noisy inputs. Specifically, we sample eight point clouds from the Stack D1 task and apply eight augmentations per sample under four noise levels:
  • Level 0: random $\mathrm{SO}(3)$ rotation only;
  • Level 1: random rotation, pointwise Gaussian jitter ($\mu=0,\;\sigma=0.05$), and random dropout, cropping, and insertion, where each of these three operations is applied to 5% of the total points;
  • Level 2: same with $\sigma=0.1$ and 10% perturbations;
  • Level 3: same with $\sigma=0.2$ and 20% perturbations.
We compare the network outputs under two conditions: (i) frozen parameters; and (ii) trained parameters. UMAP [umap] is used to project the resulting equivariant features $\mathbf{F}_{\text{equiv}}$ (shown in Fig. 3) into 2D for visualization.
As shown in Fig. 9, with frozen parameters the features remain unchanged during the training process. At Level 0, features from all eight augmentations cluster tightly for both settings, reflecting the network’s inherent $\mathrm{SO}(3)$-equivariance. As noise increases, the frozen network’s features rapidly lose their cluster structure, indicating a lack of robustness. In contrast, the trained network maintains coherent clusters at Level 1 and Level 2, showing improved tolerance to moderate perturbations. However, at Level 3 (strong jitter and 20% point violations), even the trained network fails to preserve feature consistency.
Quantitative results in Table VIII confirm these trends: without training, the estimator strictly follows equivariance but cannot handle noisy inputs; with training, it becomes more robust and reliably maps approximately rigid point clouds to a shared canonical representation, thereby improving downstream policy performance.
TABLE VIII: Ablation of freezing versus not freezing the $\mathrm{SO}(3)$-equivariant network on canonical policy performance across tasks.
Frozen Stack D1 Mug. D1 Ham. D1 Coffee D2
CP-SO3 $\times$ $68\pm 4$ $37\pm 5$ $68\pm 5$ $40\pm 7$
$18\pm 5$ $17\pm 6$ $31\pm 7$ $3\pm 2$
CP-SO2 $\times$ $76\pm 5$ $44\pm 7$ $67\pm 8$ $52\pm 8$
$68\pm 6$ $37\pm 6$ $60\pm 7$ $37\pm 6$
Figure 9: Visualization of equivariant features produced by the $\mathrm{SO}(3)$-equivariant network under different noise levels and training stages.
@@ VI Real Robot Experiments
In this section, we evaluate canonical policy on two real-world platforms: the Franka Emika Panda and UR5. We describe the experimental setup, compare against both image-based and point cloud-based baselines, and assess generalization under appearance and viewpoint shifts, as well as data efficiency in low-data regimes, followed by an analysis of failure case.
@@ VI-A Real-World Experimental Setup
Franka Emika Panda: The first system consists of a Franka Emika Panda robot arm [panda] using a relative control (see Section IV-B), equipped with a pair of Fin-Ray fingers [finray] and a fixed Intel RealSense D415 [realsense] RGB-D camera. Human demonstrations are collected via a 6-DoF 3DConnexion SpaceMouse [spacemouse], with observations and actions recorded at 10 Hz. During policy rollout, we apply DDIM [DDIM] sampling with 20 denoising steps for faster inference.
@@
We visualize the extracted point clouds on the right side of Fig. 11. RGB-D frames are converted into 3D point clouds using calibrated camera intrinsics, followed by spatial cropping within a bounding box of size $0.95\times 0.50\times 0.74~\mathrm{m}^{3}$. To enhance generalization, only the $\mathit{xyz}$ coordinates are retained, discarding color information. Consistent with the simulation setup, all point clouds are uniformly downsampled to 256 points to reduce memory usage and improve inference efficiency. For each task, we collect 100 human demonstrations with randomized object placements.
@@
We compare our method with the representative point cloud baseline DP3 [DP3], and use CP-SO2 as our canonical policy variant due to the upright alignment of the observation $z$-axis. To assess performance across modalities, we additionally evaluate two state-of-the-art image-based baselines: Diffusion Policy (DP) [diffusion_policy] and the equivariant variant EquiDiff [equidiff]. Input images are resized from $640\times 480$ to $256\times 256$, then center-cropped to $230\times 230$ before being fed into the encoder.
UR5: To test under different control and sensing conditions, we also conduct experiments on a UR5 [ur5] robot using absolute control (see Section IV-B). As shown in Fig. 12, the system is equipped with a WSG 50 gripper [wsg50] for precise width control and Fin-Ray fingers for compliant grasping. A single RealSense D415 is mounted on a mobile workstation to provide RGB-D input. The mobility of the camera facilitates easy variation of viewpoints. Demonstrations are collected via the SpaceMouse. We follow the same point cloud generation and cropping procedure as described earlier. In addition, we evaluate ACT [zhao2023learning] as an image-based baseline on the UR5, given that its chunked absolute end-effector outputs are compatible with the UR5 control mode.
On the UR5 platform, we conduct experiments to evaluate the data efficiency and robustness to camera viewpoint changes of canonical policy, comparing it against both image and point cloud policies, and additionally analyzing failure cases.
@@ VI-B Benchmarking Across Diverse Tasks
TABLE IX: Success rates over 10 trials for real-world tasks on the Franka Emika platform.
Modality Policy
Block
Stacking
Shoe
Alignment
Can
Insertion
Table
Organization
@@
[…]
Table IX summarizes the quantitative performance of different policies with various input modalities across four real-world tasks, evaluated over 10 trials on the Franka platform. CP-SO2 consistently achieves the highest success rates compared to other policies as task difficulty increases. For simpler tasks such as Block Stacking and Shoe Alignment, CP-SO2 achieves 70% and 60% success rates, outperforming the second-best policies by 40% and 10%, respectively. Besides, in these two tasks, policies based on point cloud modalities achieve higher success rates compared to those relying on image modalities.
@@
Fig. 13 illustrates the variability of policy rollouts over 10 trials. We sum the initial state observation of each trial for CP-SO2 and compute the average, which confirms that each trial differs from the others, ensuring sufficient randomness for fair evaluation. Therefore, the quantitative results reported in Table IX validate the effectiveness of the proposed canonical representation, which outperforms baselines across both point cloud and image modalities in real-world scenarios.
@@ VI-C Generalization to Unseen Object Appearances
TABLE X: Success rates over 10 trials under color variations on the Franka Emika platform.
Modality Policy
Block
Stacking
Shoe
Alignment
Can
Insertion
Table
Organization
@@
[…]
TABLE XI: Success rates over 10 trials under both color and shape variations on the Franka Emika platform.
@@
[…]
TABLE XII: Success rates over 10 trials under varying numbers of demonstrations and shoe colors on the UR5 platform.
@@
[…]
TABLE XIII: Success rates over 10 trials under varying camera viewpoints and shoe colors on the UR5 platform.
@@
In this section, to thoroughly evaluate the generalizability of canonical policy operating on point cloud inputs, we modify each task by altering object colors and shapes. The same trained policies for each task are then used to perform rollouts in the modified environments. The appearance changes are illustrated in Fig. 14.
@@
Specifically, in the Block Stacking task, both the I-shaped and T-shaped blocks are recolored to blue, representing a significant color shift from their original appearance. In the Shoe Alignment task, the shoes used during training are replaced with white shoes. Furthermore, two additional shoe types: leather shoes and hiking boots are introduced, differing substantially in both color and shape from the original shoes. In the Can Insertion task, the original black holder and orange can are replaced with a white holder and green can, respectively. Finally, in the Table Organization task, only the color of the small block is changed from black to white. Each new scenario is evaluated over 10 trials, with randomness controlled in the same manner as in Fig. 13.
@@
The experimental results under color variations are summarized in Table X. Since point cloud-based policies are trained without color information, their performance under appearance changes is significantly more robust compared to image-based policies, as evidenced by the comparison with Table IX. Notably, CP-SO2 consistently achieves the highest success rates across all tasks, outperforming the second-best policy by 30%, 30%, 10%, and 20% for tasks of increasing difficulty. Interestingly, Equidiff achieves the second-best performance in the Can Insertion task, outperforming DP3 by 20%, but shows zero success rates in other tasks, similar to DP. This phenomenon indicates that while Equidiff exhibits some degree of generalization, it remains inferior to canonical policy, as CP-SO2 maintains stable and high performance across all tasks under appearance changes.
@@
Table XI further assesses the generalization capability of each policy under simultaneous changes in both color and shape. CP-SO2 consistently achieves the highest success rates, outperforming the second-best policy by 20% in the Leather Shoe Alignment task. In the Hiking Boot Alignment task, DP3 achieves performance comparable to CP-SO2. This observation can be explained by the fact that, in the original training set, the shoes were low-top, leading the end-effector to learn grasping motions at relatively low heights. When encountering high-top hiking boots, the end-effector often follows a lower approach trajectory and collides with the boot during the grasping attempt, significantly increasing the difficulty of successful grasping. Under such challenging conditions, the performance of point cloud-based policies remains similar.
VI-D Data Efficiency Analysis
@@
In addition to evaluating policy performance under nominal conditions, we assess the generalization capability of canonical policy in real-world settings with limited data and varying observation conditions. Specifically, we study two challenging scenarios on the UR5 platform: (i) data efficiency under varying numbers of demonstrations, and (ii) robustness to camera viewpoint changes (discussed in the next subsection).
@@
Table XII summarizes the results under varying numbers of training demonstrations, comparing each method using 50 and 100 human demonstrations collected with black shoes. Each configuration is evaluated over 10 rollout trials. When trained with only 50 demonstrations, CP-SO2 is the only method that achieves non-zero success rates, reaching 40% on black shoes and 60% on white shoes. In contrast, all other baselines, including the image-based policies ACT, DP and Equidiff as well as the point cloud-based baseline DP3, fail completely under this low-data setting.
@@
As the number of demonstrations increases to 100, CP-SO2’s performance further improves to 60% and 70% on black and white shoes, respectively, outperforming the second-best method DP3 by 30% and 50%. Notably, while CP-SO2 and DP3 exhibit clear improvements with additional training data, the performance of image-based policies remains essentially unchanged.
@@
As shown in Tables IX and XII, all baselines (DP, DP3, and EquiDiff) perform worse on the UR5 compared to the Franka Panda. We attribute this to differences in the low-level control setups. In the Franka Panda experiments, we utilized an Operational Space Controller, which helped smooth out the jerky actions produced by the learned policies. In contrast, the UR5 setup relied on high-precision pose tracking without such smoothing. As illustrated in our supplementary video, under a single-camera setup and with only 50 or 100 demonstrations, baseline policies frequently produced unstable motions on the UR5, leading to degraded task performance. These results further highlight the consistent and robust performance of canonical policy across different robotic platforms.
@@
An interesting observation is that CP-SO2 achieves higher success rates on white shoes, despite being trained exclusively on black shoes, underscoring the strong generalization capability of canonical policy. From a point cloud perspective, black and white shoes share the same geometric structure, enabling a policy trained on black shoes to naturally generalize to white ones. Nevertheless, the point clouds of white shoes introduce subtle variations compared to black shoes in the training set. We believe that this slight randomness in observations may actually enhance policy robustness, a phenomenon frequently observed in imitation learning, potentially explaining the improved performance on white shoes. Introducing controlled randomness into observations is a common strategy to improve robustness of policies, as seen in techniques such as random image cropping in Diffusion Policy [diffusion_policy] and color jitter or random rotations in UMI training [chi2024universal].
VI-E Robustness to Viewpoint Shifts and Failure Case Analysis
@@
To evaluate the robustness of canonical policy under camera viewpoint changes, we fix the number of training demonstrations at 50 and rotate the mobile workstation to simulate different viewpoints. As shown in Fig. 15, the leftmost two columns (Left-Angled View) correspond to the original training setup. The middle columns (Frontal View) simulate a moderate $15^{\circ}$ rotation, while the rightmost columns (Right-Angled View) represent a larger rotation of approximately $30^{\circ}$, which induces a stronger deviation from the training viewpoint.
@@
A change in camera viewpoint is theoretically equivalent to an inverse transformation of the scene while keeping the camera fixed. This experimental setup thus enables us to validate the theoretical foundation of our canonical representation. As the mobile workstation performs rigid-body translations and in-plane rotations, the $z$-axis remains consistently aligned upright, ensuring the applicability of CP-SO2, which is designed to be $\mathrm{SO}(2)$-equivariant about the vertical axis.
@@
Table XIII summarizes the success rates under different viewpoint shifts. When the viewpoint is shifted to the Frontal View, CP-SO2 maintains strong performance with success rates of 40% and 60% for black and white shoes, respectively. In contrast, all baseline policies fail completely under the same conditions. When the camera is further rotated to the Right-Angled View, the performance of CP-SO2 drops to 30% for both shoe types.
This trend mirrors the findings in Section V-E on the noise robustness of the $\mathrm{SO}(3)$-equivariant network. Moderate viewpoint shifts (Frontal View) resemble the low-to-moderate noise levels (Levels 1–2 in Fig. 9), where parameter optimization enables the network to effectively map new observations back to the canonical frame of the original setup. Consequently, the canonicalization remains consistent and the downstream policy achieves comparable performance.
However, larger viewpoint shifts (Right-Angled View) introduce more severe disruptions: significant occlusion, previously unseen regions, and larger changes in visible geometry, which are similar to the high-noise condition (Level 3 in Fig. 9). Under such conditions, the network can no longer reliably align these altered point clouds to a single canonical space. As a result, observations that should be mapped to approximately the same canonical representation are instead projected into inconsistent frames. This increased input diversity leads to degraded policy performance.